OpenStudy (anonymous):

2x-2y=-14 what is the slope draw it

OpenStudy (p0sitr0n):

the slope is the coefficient of x in the standard form y=mx+b -2y=-14-2x y=x+7 the coefficient of x is one

OpenStudy (p0sitr0n):



OpenStudy (anonymous):

thanks